# Lie algebra homology with coefficients tensor products of the adjoint representation in relative polynomial degree $2$

### v1

@@ Abstract.
There is a splitting of this homology by polynomial degree (for polynomiality with respect to the generators of the free Lie algebra) and one can consider the polynomial degree relative to the number of tensor factors in the coefficients. It suffices to consider the Lie algebra homology in homological degree one; this vanishes in relative degree 00 and is readily calculated in relative degree $1$.
@@
Key words and phrases:
Lie algebra homology; adjoint representation; polynomial functor
@@ 1. Introduction
Here we are interested in the case where $\mathfrak{g}$ is the free Lie algebra $\mathbb{L}\mathrm{ie}(V)$ on a finite-dimensional $\mathbb{Q}$-vector space $V$, considered as a functor of $V$. This is motivated by the relationship with higher Hochschild homology arising from the author’s joint work with Christine Vespa [PV25]. More explicitly, in that paper we established the relationship with the category of outer functors, namely functors from $\mathbf{gr}$ (the category of finitely-generated free groups) to $\mathbb{Q}$-vector spaces on which inner automorphisms of free groups act trivially. In [Pow24b], using the infinitesimal interpretation of polynomial functors on free groups given in [Pow24a], this was related to Lie algebra homology with coefficients in tensor products of the adjoint representation. This is also related to work of Nir Gadish and Louis Hainaut [GH24], who study the homology of certain configuration spaces and relate this to higher Hochschild homology; their calculations are used by Hainaut in [Hai23] to study homological properties of the category of polynomial outer functors.
Working with the free Lie algebra, the question reduces to that of calculating the representations of $\mathfrak{S}_{r}$
@@
$H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})$
naturally with respect to $V$. For $r=1$, the $H_{1}$ term is given by the Schur functor associated to the cyclic Lie operad (cf. [Lev02], for example):
@@
One can restrict to studying $H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})$, since the homological degree zero term can be derived from this. This is advantageous, as observed by Gadish and Hainaut, since $H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})$ is usually smaller than $H_{0}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})$ in a precise sense.
@@
where $\mathrm{triv}_{r}$ is the trivial representation of $\mathfrak{S}_{r}$ and $\Gamma^{r+1}(V)=(V^{\otimes r+1})^{\mathfrak{S}_{r+1}}$ is the $(r+1)$st divided power functor.
The purpose of this paper is to treat the next case, $n=r+2$. In the following statement, $\mathrm{sgn}_{r}$ denotes the sign representation of $\mathfrak{S}_{r}$, $\Lambda^{r+2}$ is the $(r+2)$nd exterior power functor and $\mathbf{S}_{(r,1^{2})}$ is the Schur functor associated to the partition $(r,1^{2})$.
@@
Theorem 1 is straightforward for $r\in\{0,1,2\}$. The case $r=3$ is established in Section 5. (This case was already known by computer calculation, for example as a consequence of calculations in [GH24]; it is established here by hand.) These cases serve as the initial steps for the inductive proof that is given in Section 6.
@@
(2) $\Phi^{1}[r+2,r]\cong S_{(r+2)}\boxtimes S_{(r)}\ \oplus\ S_{(3,1^{r-1})}\boxtimes S_{(1^{r})}$
(using this paper’s notation for the simple representations). Observe that the right hand side of (2) is obtained from (1) by replacing each occurrence of $S_{\lambda}$ for a partition $\lambda$ by $S_{\lambda^{\dagger}}$, where $\lambda^{\dagger}$ is the conjugate partition to $\lambda$.
In [GH24, Section 4.4], Gadish and Hainaut explain how their objects $\Phi[-,-]$ are related to higher Hochschild homology as considered in [PV25] and by Turchin and Willwacher in [TW19]. In particular, this explains how $\Phi^{1}[r+2,r]$ is related to $H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})^{[r+2]}$, also using the relationship between higher Hochschild homology and the Lie algebra homology considered here that is explained by the results of [PV25] in conjunction with those of [Pow24a] and [Pow24b]. In particular, as already observed in [PV25], this accounts for the twisting that results in the passage $\lambda\mapsto\lambda^{\dagger}$ observed above.
@@ 2.1. Representations of the symmetric groups and Schur functors
We work over the rational numbers, $\mathbb{Q}$. The category of finite-dimensional $\mathbb{Q}$-vector spaces is denoted by $\mathbf{mod}_{\mathbb{Q}}$; all undecorated tensor products $\otimes$ are understood to be taken over $\mathbb{Q}$. For $n\in\mathbb{N}$, the symmetric group on $n$ letters is denoted $\mathfrak{S}_{n}$. The category of $\mathbb{Q}\mathfrak{S}_{n}$-modules (or $\mathfrak{S}_{n}$-modules) is semi-simple. The simple $\mathfrak{S}_{n}$-modules are indexed by the partitions $\lambda$ of $n$, with $S_{\lambda}$ denoting the associated simple module. The indexing convention is chosen so that $S_{(n)}$ is the trivial representation $\mathrm{triv}_{n}$ and $S_{(1^{n})}$ the sign representation $\mathrm{sgn}_{n}$. For groups $H\subset G$, the notation $\downarrow^{G}_{H}$ denotes restriction from $G$-modules to $H$-modules and $\uparrow_{H}^{G}$ denotes induction.
@@
This is a polynomial functor that is homogeneous of degree $n$. (See [Mac15, Appendix I.A] for polynomial functors in characteristic zero.) The Schur functor construction establishes an equivalence of categories between $\mathfrak{S}_{n}$-modules and homogeneous polynomial functors of degree $n$; given such a polynomial functor $F$, the associated $\mathfrak{S}_{n}$-representation is given by the natural transformations
$\mathrm{Nat}(T^{n},F),$
@@
For $\lambda=(n)$, the Schur functor $\mathbf{S}_{(n)}$ identifies as the $n$th symmetric power functor $V\mapsto S^{n}(V)=(V^{\otimes n})/\mathfrak{S}_{n}$. Over $\mathbb{Q}$, this is isomorphic to the $n$th divided power functor $V\mapsto\Gamma^{n}(V)=(V^{\otimes n})^{\mathfrak{S}_{n}}$. In particular, there is a canonical surjection $V^{\otimes n}\twoheadrightarrow S^{n}(V)$ and a canonical inclusion $\Gamma^{n}(V)\hookrightarrow V^{\otimes n}$. For $\lambda=(1^{n})$, the Schur functor $\mathbf{S}_{(1^{n})}$ identifies as the $n$th exterior power functor $V\mapsto\Lambda^{n}(V)$.
@@
We record the following basic fact which plays the rôle of Schur’s lemma:
@@
For a partition $\lambda$, $l(\lambda)$ denotes its length, i.e., the number of non-zero terms in the partition. One has the following standard result on the connectivity of the simple Schur functors:
@@ 2.2. A form of the de Rham differential
For $a<b\in\mathbb{N}$, $\mathsf{d}$ is the unique map which fits into the commutative diagram of natural transformations
where the bottom map is induced by the product $\mu:\Lambda^{a}\otimes\Lambda^{1}\rightarrow\Lambda^{a+1}$ and the vertical maps correspond are induced by the canonical inclusion $\Gamma^{*}\subset T^{*}$.
As usual, one can identify the kernel and cokernel of $\mathsf{d}$ using the Pieri rule:

### v4

@@ Abstract.
There is a splitting of this homology by polynomial degree (for polynomiality with respect to the generators of the free Lie algebra) and one can consider the polynomial degree relative to the number of tensor factors in the coefficients. It suffices to consider the Lie algebra homology in homological degree one; this vanishes in relative degree 0 and is readily calculated in relative degree $1$.
@@
Keywords
Lie algebra homology; adjoint representation; polynomial functor.
@@ 1. Introduction
Here we are interested in the case where $\mathfrak{g}$ is the free Lie algebra $\mathbb{L}\mathrm{ie}(V)$ on a finite-dimensional $\mathbb{Q}$-vector space $V$, considered as a functor of $V$. This is motivated by the relationship with higher Hochschild homology arising from the author’s joint work with Christine Vespa [PV25]. More explicitly, in that paper we established the relationship with the category of outer functors, namely functors from $\mathbf{gr}$ (the category of finitely-generated free groups) to $\mathbb{Q}$-vector spaces on which inner automorphisms of free groups act trivially. In [Pow24b], using the infinitesimal interpretation of polynomial functors on free groups given in [Pow24a], this was related to Lie algebra homology with coefficients in tensor products of the adjoint representation. This is also related to work of Nir Gadish and Louis Hainaut [GH24], who study the homology of certain configuration spaces and relate this to higher Hochschild homology; their calculations are used by Hainaut in [Hai23] to study homological properties of the category of polynomial outer functors. (See Appendix A, which is independent of the body of the text, for slightly more detail on the above.)
Working with the free Lie algebra, since $H_{*}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})$ vanishes for >1*>1, its calculation reduces to that of the $\mathfrak{S}_{r}$-modules
@@
$H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r}),$
naturally with respect to $V$. For $r=1$, the $H_{1}$-term is given by the Schur functor associated to the cyclic Lie operad (see Example 3.1):
@@
One can restrict to studying $H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})$, since the homological degree zero term can be related to this by using the isomorphism
$H_{0}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})\ \oplus\ (V\otimes\mathbb{L}\mathrm{ie}(V)^{\otimes r})\ \cong\ H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})\ \oplus\ \mathbb{L}\mathrm{ie}(V)^{\otimes r}$
derived from the complex calculating the homology. This is advantageous, as observed by Gadish and Hainaut, since $H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})$ is usually smaller than $H_{0}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})$.
@@
where $\mathrm{triv}_{r}$ is the trivial representation of $\mathfrak{S}_{r}$ and $\Gamma^{r+1}(V)=(V^{\otimes r+1})^{\mathfrak{S}_{r+1}}$ is the (r+1)(r+1)st divided power functor.
The purpose of this paper is to treat the next case, $n=r+2$. In the following statement, $\mathrm{sgn}_{r}$ denotes the sign representation of $\mathfrak{S}_{r}$, $\Lambda^{r+2}$ is the (r+2)(r+2)nd exterior power functor and $\mathbf{S}_{(r,1^{2})}$ is the Schur functor associated to the partition $(r,1^{2})$.
@@
Here, for $\lambda$ a partition of $n\in\mathbb{N}$, $S_{\lambda}$ denotes the corresponding simple representation of $\mathfrak{S}_{n}$. With our conventions, $S_{(n)}$ is isomorphic to $\mathrm{triv}_{n}$ and $S_{(1^{n})}$ to $\mathrm{sgn}_{n}$.
Theorem 1 is straightforward for $r\in\{0,1,2\}$. The case $r=3$ is a technical calculation that is postponed until Section 6. (This case was already known by computer calculation, for example as a consequence of calculations in [GH24]; here it is established by hand.) These cases serve as the initial steps for the inductive proof that is given in Section 5.
@@
(2) $\Phi^{1}[r+2,r]\cong S_{(r+2)}\boxtimes S_{(r)}\ \oplus\ S_{(3,1^{r-1})}\boxtimes S_{(1^{r})}.$
Observe that the right hand side of (2) is obtained from (1) by replacing each occurrence of $S_{\lambda}$ for a partition $\lambda$ by $S_{\lambda^{\dagger}}$, where $\lambda^{\dagger}$ is the conjugate partition to $\lambda$.
In [GH24, Section 4.4], Gadish and Hainaut explain how their objects $\Phi[-,-]$ are related to higher Hochschild homology, as considered in [PV25] and by Turchin and Willwacher in [TW19]. Now, [PV25, Theorem 6] gives a description of higher Hochschild homology in terms of algebraic objects defined in terms of functors on the category $\mathbf{gr}$. (This result also explains the passage between $\lambda$ and $\lambda^{\dagger}$ for the action of $\mathfrak{S}_{r+2}$ in the case at hand.)
These algebraic objects then relate to the Lie algebra homology using the results of [Pow24a] and [Pow24b] (see Appendix A for an outline of the main ideas). Indeed, the algebraic objects appearing in [PV25, Theorem 6] are essentially determined by the family of functors $\omega\beta_{s}\mathbb{Q}\mathfrak{S}_{s}$, for $s\in\mathbb{N}$, that appear in [PV25, Theorem 7]. By that theorem, these form a family of injective cogenerators in the appropriate category of outer polynomial functors on $\mathbf{gr}$. After dualizing (so as to pass to functors on $\mathbf{gr}^{\mathrm{op}}$), this corresponds to the property exhibited in Example A.7, thereby establishing the link with Lie algebra homology.
Putting the above together yields the explicit relationship between $\Phi^{1}[r+2,r]$ and $H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})^{[r+2]}$ which shows that Theorem 1 implies Corollary 1.
@@ Acknowledgement
The author thanks all the anonymous referees for their careful reading of the paper and, in particular, for their valuable suggestions, which have lead to an improvement of the exposition.
@@ 2.1. Representations of the symmetric groups and Schur functors
We work over the rational numbers, $\mathbb{Q}$. The category of finite-dimensional $\mathbb{Q}$-vector spaces is denoted by $\mathbf{mod}_{\mathbb{Q}}$; all undecorated tensor products $\otimes$ are understood to be taken over $\mathbb{Q}$. For $n\in\mathbb{N}$, the symmetric group on $n$ letters is denoted $\mathfrak{S}_{n}$. The category of $\mathbb{Q}\mathfrak{S}_{n}$-modules (or $\mathfrak{S}_{n}$-modules) is semisimple. The isomorphism classes of simple $\mathfrak{S}_{n}$-modules are indexed by the partitions $\lambda$ of $n$ (written $\lambda\vdash n$), with $S_{\lambda}$ denoting the associated simple module. The indexing convention is chosen so that $S_{(n)}$ is the trivial representation $\mathrm{triv}_{n}$ and $S_{(1^{n})}$ the sign representation $\mathrm{sgn}_{n}$. For groups $H\subset G$, the notation $\downarrow^{G}_{H}$ denotes restriction from $G$-modules to $H$-modules and $\uparrow_{H}^{G}$ denotes induction.
@@
This is a polynomial functor that is homogeneous of degree $n$. (See [Mac15, Appendix I.A] for polynomial functors in characteristic zero.) The Schur functor construction establishes an equivalence of categories between $\mathfrak{S}_{n}$-modules and homogeneous polynomial functors of degree $n$; given such a polynomial functor $G$, the associated $\mathfrak{S}_{n}$-representation is given by the natural transformations
$\mathrm{Nat}(T^{n},G),$
@@
There is more structure: given two polynomial functors $F$ and $G$, one can form their pointwise tensor product, defined by $(F\otimes G)(V):=F(V)\otimes G(V)$. If $F$ and $G$ are homogeneous polynomial of degrees $m$ and $n$ respectively, then $F\otimes G$ is homogeneous polynomial of degree $m+n$. At the level of representations of the symmetric groups, this corresponds to the functor which sends a $\mathbb{Q}\mathfrak{S}_{m}$-module $M$ and a $\mathbb{Q}\mathfrak{S}_{n}$-module $N$ to the $\mathbb{Q}\mathfrak{S}_{m+n}$-module $(M\boxtimes N)\uparrow_{\mathfrak{S}_{m}\times\mathfrak{S}_{n}}^{\mathfrak{S}_{m+n}}$.
@@
For $\lambda=(n)$, the Schur functor $\mathbf{S}_{(n)}$ identifies as the $n$th symmetric power functor $V\mapsto S^{n}(V)=(V^{\otimes n})/\mathfrak{S}_{n}$. Over $\mathbb{Q}$, this is isomorphic to the $n$th divided power functor $V\mapsto\Gamma^{n}(V)=(V^{\otimes n})^{\mathfrak{S}_{n}}$. There is a canonical surjection $V^{\otimes n}\twoheadrightarrow S^{n}(V)$ and a canonical inclusion $\Gamma^{n}(V)\hookrightarrow V^{\otimes n}$. For $\lambda=(1^{n})$, the Schur functor $\mathbf{S}_{(1^{n})}$ identifies as the $n$th exterior power functor $V\mapsto\Lambda^{n}(V)$.
@@
To make these structures more explicit, we consider the evaluation on $V$ a vector space of dimension $t$, with chosen basis $\{x_{1},\ldots,x_{t}\}$. Then $S^{*}(V)$ identifies as the polynomial algebra $\mathbb{Q}[x_{1},\ldots,x_{t}]$; $S^{d}(V)$ corresponds to the subspace of homogeneous polynomials of degree $d$. Then the products $S^{a}\otimes S^{b}\rightarrow S^{a+b}$ (evaluated on $V$) are given by the product in this polynomial algebra. The coproducts (evaluated on $V$) are given by equipping the polynomial algebra with the primitively-generated Hopf algebra structure with coproduct $\Delta$ determined by $\Delta x_{i}=x_{1}\otimes 1+1\otimes x_{i}$. Thus, for example, the coproduct $S^{2}\rightarrow S^{1}\otimes S^{1}$ is determined by $vw\mapsto v\otimes w+w\otimes v$, for any $v,w\in V$.
The exterior power functors can be treated by the same argument, by introducing a grading and using Koszul signs. Namely, we form the free graded commutative algebra on $sV$ (where $sV$ is $V$ placed in degree one) and equip this with the primitively-generated Hopf algebra structure. The underlying object can be identified with the exterior algebra on $x_{1},\ldots,x_{t}$, where the grading corresponds to the polynomial degree. The product (evaluated on $V$) then identifies as the wedge product $\wedge$. The coproduct is identified as in the case of the symmetric powers; for example, the coproduct (evaluated on $V$) $\Lambda^{2}\rightarrow\Lambda^{1}\otimes\Lambda^{1}$ is given by $v\wedge w\mapsto v\otimes w-w\otimes v$.
Example 2.3.
For two partitions $\lambda\vdash d$ and $\mu\vdash e$, the tensor product of the associated Schur functors $\mathbf{S}_{\lambda}$ and $\mathbf{S}_{\mu}$ is given by the Littlewood-Richardson rule. Namely, there is an isomorphism
$\mathbf{S}_{\lambda}\otimes\mathbf{S}_{\mu}\cong\bigoplus_{\nu\vdash d+e}\mathbf{S}_{\nu}^{\bigoplus c^{\nu}_{\lambda\mu}},$
where $c^{\nu}_{\lambda\mu}$ are the Littlewood-Richardson coefficients. Using the properties of the Schur functor construction recalled above, this follows from the classical Littlewood-Richardson rule for representations of the symmetric groups
$(S_{\lambda}\boxtimes S_{\mu})\uparrow_{\mathfrak{S}_{d}\times\mathfrak{S}_{e}}^{\mathfrak{S}_{d+e}}\cong\bigoplus_{\nu\vdash d+e}S_{\nu}^{\bigoplus c^{\nu}_{\lambda\mu}}.$
Specializing to the case where $\mu=(e)$, so that $\mathbf{S}_{\mu}$ is the symmetric power $S^{e}$, this gives the Pieri formula:
$\mathbf{S}_{\lambda}\otimes S^{e}\cong\bigoplus_{\begin{subarray}{c}\nu\vdash d+e\\ \nu/\lambda\in\mathrm{HS}\end{subarray}}\mathbf{S}_{\nu}.$
Here, the condition $\nu/\lambda\in\mathrm{HS}$ states that the skew partition $\nu/\lambda$ is a horizontal strip (using the convention that the Young diagram associated to a partition $\lambda$ has $\lambda_{i}$ boxes in the $i$th row). The horizontal strip condition means that the Young diagram of $\nu$ is obtained from that of $\lambda$ by adding at most one box in each column. (See [Wey03, Corollary 2.3.5], for example, paying attention to the fact that Weyman uses a different indexing convention.)
Taking $e=1$, this reduces to the classical Pieri rule: the sum is over all partitions $\nu\vdash d+1$ such that the Young diagram of $\nu$ contains that of $\lambda$.
There is an (equivalent) formulation of the Pieri formula for the case $\mu=(1^{e})$, so that $\mathbf{S}_{\nu}$ identifies as the exterior power functor $\Lambda^{e}$. In this case:
$\mathbf{S}_{\lambda}\otimes\Lambda^{e}\cong\bigoplus_{\begin{subarray}{c}\nu\vdash d+e\\ \nu/\lambda\in\mathrm{VS}\end{subarray}}\mathbf{S}_{\nu},$
where $\mathrm{VS}$ now refers to vertical strips (exchanging the rôle of horizontal and vertical).
We record the following basic fact, which plays the rôle of Schur’s lemma:
@@
For a partition $\lambda$, $l(\lambda)$ denotes its length, i.e., the number of non-zero terms in the partition. One has the following standard result on the connectivity of the simple Schur functors, which is a constituent of Schur-Weyl duality:
@@
Proof.
(For completeness, we sketch a proof.) There is an explicit description of the Schur functors (see, for example, [Wey03, Chapter 2], noting that Weyman works over a general commutative ring and uses a different indexing convention). This implies, in particular, that $\mathbf{S}_{\lambda}$ is a subfunctor of $\bigotimes_{j=1}^{\lambda_{1}}\Lambda^{\mu_{j}}$, where $\mu:=\lambda^{\dagger}$ is the conjugate partition, which has length $l(\mu)=\lambda_{1}$ (correspondingly, $\mu_{1}=l(\lambda)$). The vanishing statement follows from the fact, for $t\in\mathbb{N}$, $\Lambda^{t}(V)$ is zero if and only if $\dim V<t$.
That $\mathbf{S}_{\lambda}(V)$ is non-zero for $\dim V=l(\lambda)$ follows from further analysis of the construction of $\mathbf{S}_{\lambda}$. For example, this follows from [Wey03, Proposition 2.1.4]. ∎
@@ 2.2. A form of the de Rham differential
For $a,b\in\mathbb{N}$ with $b>0$, $\mathsf{d}$ is the unique map which fits into the commutative diagram of natural transformations
(8) ${\Lambda^{a}\otimes\Gamma^{b}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathsf{d}}$${\Lambda^{a+1}\otimes\Gamma^{b-1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Lambda^{a}\otimes T^{b}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Lambda^{a+1}\otimes T^{b-1},}$
where the bottom map is induced by the product $\mu:\Lambda^{a}\otimes\Lambda^{1}\rightarrow\Lambda^{a+1}$ and the vertical maps are induced by the canonical inclusion $\Gamma^{*}\subset T^{*}$.
One has the well-known identification of the kernel and cokernel of $\mathsf{d}$:
@@
Proof.
We use the Pieri formula (see Example 2.3). This implies that, for positive integers $a$ and $b$, there is an isomorphism
$\Lambda^{a}\otimes\Gamma^{b}\cong\mathbf{S}_{(b+1,1^{a-1})}\oplus\mathbf{S}_{(b,1^{a})}.$
The differential $\mathsf{d}:\Lambda^{a}\otimes\Gamma^{b}\rightarrow\Lambda^{a+1}\otimes\Gamma^{b-1}$ is clearly non-zero and, by the above, identifies as a natural transformation
$\mathbf{S}_{(b+1,1^{a-1})}\oplus\mathbf{S}_{(b,1^{a})}\stackrel{{\scriptstyle\mathsf{d}}}{{\rightarrow}}\mathbf{S}_{(b,1^{a})}\oplus\mathbf{S}_{(b-1,1^{a+1})}.$
The result thus follows from Schur’s lemma (Proposition 2.4 here). ∎
@@
Hence, by Lemma 2.8, one has
$\mathbf{S}_{(r,1^{2})}\cong\ker\big(\Lambda^{3}\otimes\Gamma^{r-1}\hookrightarrow\Lambda^{3}\otimes T^{r-1}\cong\Lambda^{3}\otimes\Lambda^{1}\otimes T^{r-2}\rightarrow\Lambda^{4}\otimes T^{r-2}\big),$
where the first map is given by the inclusion $\Gamma^{r-1}\subset T^{r-1}$ and the second by the product $\Lambda^{3}\otimes\Lambda^{1}\rightarrow\Lambda^{4}$, as in that Lemma.